Tohban Report 2016-11-30
Tohban Reports | |
---|---|
Start Date: | 16 November 2016 |
End Date: | 30 November 2016 |
Tohban: | Pascal Saint-Hilaire |
Tohban email: | Pascal@ssl.berkeley.edu |
Next Tohban: | Juan Carlos Martinez Oliveros |
List all reports |
Solar Activity
B-flare activity How many GOES flares occurred?
Flares above B, C, M, X class were 21 1 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 10 0 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 7 0 0 0
There were RHESSI flares/GOES flares 29 / 22 over the time range 22-Nov-16 29-Nov-16
Activity has picked up in the last couple of days. Two ~M1 flare peaks missed by RHESSI.
Memory Management
Decimation was normal/vigorous for the full time period. The SSR was at 28.6% with four passes still to go. Detectors 3 and 9 are on full-time, including spacecraft night, and rear segment decimation was on in high-latitude regions.
Spacecraft Status
Detectors 3 and 9 are recording events.
Data Gaps
No gaps (not even 300 s ones) before November 29th.
DATA GAPS FOR 2016/11/29
TIME RANGE: 2016-11-29T00:00:00.000 -- 2016-11-30T00:00:00.000 NO GAPS IN APP_ID = 1 (VC1-SOH) NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2016-11-29T19:20:00.000 -- 2016-11-30T00:00:00.000 16800.000
Detector issues
Detectors 3 and 9 are powered on, and are recording events. Some threshold changes:
2016-11-28T18:30:02 D9R fast thresholds from 0xD0 to 0xE0.
2016-11-28T21:45:07 D9R fast thresholds from 0xE0 to 0xFF.
2016-11-28T21:45:33 D3R fast thresholds from 0xA0 to 0xB0.
Notes
No quicklook image for the 2016-11-28T07:00 C7.5 flare. Jim & Albert looking into this.
Spacecraft Management
Decimation | Normal/Vigorous |
HLAT Decimation | Rear decimation |
Night time data (fronts) | full-time |
Night time data (rears) | full-time |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |